# Cost-Effective, Low Latency Vector Search with Azure Cosmos DB

### v1

@@ 4.4. Recall Stability over Updates
Figure 11. Recall trends for WikiCohere-10M expiration time runbook. Vectors are compressed to 192 bytes, and the index uses $R=32,L_{build}=100,L_{search}=200$.
[…]
[Recall trends for WikiCohere-10M expiration time runbook]Recall trends for WikiCohere-10M expiration time runbook
Figure 12. Recall trends for the MSTuring-1M expiration time runbook. Vectors are compressed to 50 bytes, and the index uses $R=32,L_{build}=100,L_{search}=100$.
[…]
[Recall trends for the MSTuring-1M expiration time runbook]Recall trends for the MSTuring-1M expiration time runbook
Figure 13. Recall trends for the MSTuring-10M clustered runbook. Vectors are compressed to 50 bytes, and the index uses $R=32,L_{build}=100,L_{search}=100$.
[…]
[Recall trends for the MSTuring-10M clustered runbook]Recall trends for the MSTuring-10M clustered runbook
We now consider the recall stability of a vector index in Cosmos DB using the runbooks described at the beginning of this section, plotted in Figures 1112, and 13. In all runbooks, an initial quantization is computed after 1000 vectors are inserted, and the index is re-quantized after 25000 vectors are inserted. The experiments compare the effects of taking no action other than dropping the deleted vector (labeled ”Without_Inplace_Delete”) versus using the inplace delete algorithm (Algorithm 6).
The experiments show that the in-place delete algorithm is critical for maintaining high recall under a stream of updates. In the case of the expiration time runbooks (Figures 11 and 12), in-place delete increases average recall by 1-3 percentage points. On the clustered runbook, which experiences distribution shift and is significantly more difficult than the expiration time runbooks, in-place deletes achieve recall that is as much as 20 percentage points higher than the baseline, meaning they are significantly more robust to distribution shifts.
@@ 4.5. Sharded indices for Multi-tenancy
Table 3. Sharded Diskann with year as the shard key (row 1) compared with non-sharded DiskANN (rows 2 and 3) on YFCC.
@@ 5.1. Algorithms for Vector Search
Graph-based algorithms form a proximity graph with one vertex per embedding and greedily traverse the graph at query time to find a query’s nearest neighbors. Examples include HNSW (Malkov and Yashunin, 2016), NSG (Fu et al., 2019), DiskANN (Subramanya et al., 2019), and ELPIS (Azizi et al., 2023). Empirically, the query complexity of graphs scales logarithmically with the size of the dataset and is much better than LSH and clustering (see Figure 14). Graph-based ANNS algorithms can maintain high recall in the streaming setting (Singh et al., 2021; Xu et al., 2025b). Furthermore, they are adaptable to filtered queries (Gollapudi et al., 2023) and queries that respect notions of diversity (Anand et al., 2025). For these reasons—scalability, query efficiency, versatility, and robust updates—we use graphs in this work.
@@ 5.2. Vector Indices in Databases
JVector (Jonathan Ellis et al., [n.d.]) is a Cassandra based vector index with DiskANN. They construct a DiskANN index in one shot and ingest it into the database, and support incremental insertions and deletions using FreshDiskANN. Details on how the indices are managed and updated across partitions and segments are not documented to the best of our knowledge. DataStax offers a managed product based on JVector. While their serverless offering does not provide an availability SLA, the enterprise provisioned capacity model has a 99.99% SLA. While their enterprise pricing is unspecified, their standard tier has a monthly price of \$900/month for 10 Queries/second over a million sized index. The corresponding Cosmos DB monthly cost with Autoscale is less than \$50.
Elastic offers vector search using a segment-based indexing system. It therefore has disadvantages similar to the SingleStore-V design. Furthermore, as a managed search engine, it does not offer the same level of robustness and data consistency as as an operational database. To compare costs, consider the 10 million vector index over Wiki-Cohere data. Even with scalar quantization, we would need a machine with 15GB DRAM. We estimate using their pricing tool that the smallest such machine from Vector Search Optimized SKUs with 1.9vCPUS would costs about \$0.7/hour or about \$500/month. A manually provisioned Cosmos DB service for 500 RU/sec or 5000 RU/sec would serve 10 QPS and 100 QPS and cost about \$$\$30$ and \$$\$300$/month respectively.
pgvector (Andrew Kane et al., [n.d.]) integrates HNSW and IVF-Flat indexing algorithm inside PostgreSQL and offers the inherent advantages of the PosgreSQL ecosystem. Initial versions required machines large enough to fit the entire dataset as well as the index in memory for reasonable performance. Scalar and binary quantizations have recently been added to reduce the memory requirement, but performant indexing and updates still rely on the availability of much larger memory than our system. pgvectorscale (Matvey Arye et al., [n.d.]) similarly integrates DiskANN inside PostgreSQL and offers better performance than pgvector in limited memory settings. We leave an exhaustive benchmarking against these systems for future work. They primarily differ from our system in that they do not offer flexible schemas, multi-tenancy, and scale-out out of the box. V-Base (Zhang et al., 2023) also builds on PostgreSQL and aims to improve vector search performance. However it uses an independent buffer pool not integrated with to PostgreSQL storage engines. Substantially more work is needed to build high-availability and other features in this system.
@@ 5.3. Specialized Serverless Vector DBs
For example, Pinecone offers 99.95% availability, but maintains only one data replica by default and provides limited backup-and-restore functionality applicable solely to the vector index, not the data itself. In terms of security, Pinecone supports Role-Based Access Control (RBAC) for both control and data plane operations without custom role support. Certifications include HIPAA BAA, AICPA SOC, GDPR, and ISO 27001 as of this writing.
Another offering, Zilliz, provides no availability SLA, lacks replication and backup-and-restore, and limits RBAC functionality to the control plane without custom roles. Its compliance certifications are limited to SOC 2 Type II, ISO/IEC 27001, and GDPR.
Turbopuffer (Eskildsen, 2024) is another serverless vector database designed for multi-tenant usage. However, it still lacks significant enterprise and robustness features such as RBAC, backup-and-restore, global availability. Turbopuffer does not publicly list query price for 10 million vector shards, however, its query complexity increases substantially with the size of the tenant. For example, cost per 1 million queries increases from \$3.58 to \$33.4 when the number of documents per name space increases from 100,000 to a modest 1 million. (Turbopuffer, 2025)
@@ 7. Acknowledgements
We thank Phil Bernstein for his feedback on the work and on improving this draft.
@@ Appendix A DiskANN plots and algorithms
      return [[[[closest $k$ points from $\mathcal{V}$ with label $\mathcal{B}$; $\mathcal{V}$]]]]
@@
Note that in order to use Algorithm 8, the product quantization must produce distances that can be compared with full-precision distances (i.e. they must preserve approximate distances, not merely an approximate ordering).
Data: Graph $G$, point $p\in P$, candidate set $\mathcal{E}$, distance threshold $\alpha\geq 1$, degree bound $R$, maximum full-precision vector allowance $M$. Functions FullPrecisionDist, FullQuantDist, QuantQuantDist indicate distance calculations between two full-precision vectors, a full-precision vector and a quantized vector, and two quantized vectors, respectively.
Result: $G$ is modified by setting at most $R$ new out-neighbors for $p$ using at most $M$ full-precision vector lookups
begin
      $\mathcal{E}\leftarrow(\mathcal{E}\cup N_{\rm out}({p}))\setminus\{p\}$
       // sort $\mathcal{E}$ by quantized distance from $p$ and retain distances
       $\mathcal{E}\leftarrow$ sort $\mathcal{E}$ by quantized distance from ${\sf x}_{p}$
       // overwrite closest $M$ nodes with full distances, then sort again
       for $i,({\sf x}_{q},dist)\in$ $\mathcal{E}$) do
             if $i<M$ then
                   $dist\leftarrow$ FullPrecisionDist(${\sf x}_{p},{\sf x}_{q}$)
      $\mathcal{E}\leftarrow$ sort $\mathcal{E}$ by recorded distance from ${\sf x}_{p}$
      $\mathcal{N}\leftarrow\emptyset$
       for $i,({\sf x}_{q},dist\_qp)\in\mathcal{E}$ do
             for $(j,({\sf x}_{r},dist\_rp))\in N$ do
                   $dist\_qr\leftarrow 0$
                   $add\leftarrow\text{True}$
                   if $i<M\&j<M$ then
                         $dist\_qr\leftarrow$ FullPrecisionDist(${\sf x}_{r},{\sf x}_{q}$)
                  else if $i\geq M\&j<M$ then
                         $dist\_qr\leftarrow$ FullQuantDist(${\sf x}_{r},{\sf x}_{q}$)
                  else
                         $dist\_qr\leftarrow$ QuantQuantDist(${\sf x}_{r},{\sf x}_{q}$)
                  if $\alpha*dist\_rp<dist\_rq$ then
                         $add\leftarrow\text{False}$
                         break
            if add then
                   $\mathcal{N}\leftarrow\mathcal{N}\cup\{(i,({\sf x}_{q},dist\_qp)\}$
            if $\mathcal{N}|=R$ then
                   break
      $N_{\rm out}({p})\leftarrow\{q\text{ for }(i,({\sf x}_{q},dist\_qp))\in\mathcal{N}\}$
Algorithm 8 ${\rm QuantizedRobustPrune}({p},{\mathcal{E}},{\alpha},{R}{M})$
@@ Appendix B Cosmos DB and DiskANN interoperability
To insert a vector, the Cosmos DB IndexManager calls diskann.insert with appropriate execution context, which in turn accesses the Bw-Tree index and the Document Store. The DiskANN library written in Rust is sandwiched between IndexManager and Bw-Tree/Document Store components written in C++. The interleaving is similar for the query path, and is illustrated in Figure 15. Since all of these components are written in an asynchronous programming style with different languages and run times, this leads to an interesting challenge of implementing asynchronous callbacks across language boundary. The standard way to integrate Rust/C++ using C ABI has no native support for async calls, and public FFI libraries that support C++/Rust async inter-op are non-existent. We build an internal solution to address this, and did not observe any significant overheads from this interface.

### v2

@@ 4.5. Recall Stability over Updates
(a) WikiCohere-10M expiration time runbook.
(b) MS Turing-1M expiration time runbook.
(c) MSTuring-10M clustered runbook.
Figure 13. Recall trends for the runbooks described in Section 4. All runs used index parameters $R=32,L_{build}=100,L_{search}=200$. WikiCohere was compressed to 192 bytes and MSTuring to 50 bytes.
We now consider the recall stability of a vector index in Cosmos DB using the runbooks described at the beginning of this section, plotted in Figures 13(a)13(b), and 13(c). In all runbooks, an initial quantization is computed after 1000 vectors are inserted, and the index is re-quantized after 25000 vectors are inserted. The experiments compare the effects of taking no action other than dropping the deleted vector (labeled ”Drop Policy” in the plots) versus using the inplace delete algorithm (Algorithm 6).
The experiments show that the in-place delete algorithm is critical for maintaining high recall under a stream of updates. In the case of the expiration time runbooks (Figures 13(a) and 13(b)), in-place delete increases average recall by 1-3 percentage points. On the clustered runbook, which simulates distribution shift and is significantly more difficult than the expiration time runbooks, in-place deletes achieve recall that is as much as 20 percentage points higher than the baseline, meaning they are significantly more robust to distribution shifts.
@@ 4.6. Sharded indices for Multi-tenancy
Table 3. Sharded Diskann with year as the shard key (row 1) compared with non-sharded DiskANN (rows 2, 3) on YFCC.
@@ 5.1. Algorithms for Vector Search
Graph-based algorithms form a proximity graph with one vertex per embedding and greedily traverse the graph at query time to find a query’s nearest neighbors. Examples include HNSW (Malkov and Yashunin, 2016), NSG (Fu et al., 2019), DiskANN (Subramanya et al., 2019), and ELPIS (Azizi et al., 2023). Empirically, the query complexity of graphs scales logarithmically with the size of the dataset and is much better than LSH and clustering (see Figure 14). Graph-based ANNS algorithms can maintain high recall in the streaming setting (Singh et al., 2021; Xu et al., 2025). Furthermore, they are adaptable to filtered queries (Gollapudi et al., 2023) and queries that respect notions of diversity (Anand et al., 2025). For these reasons—scalability, query efficiency, versatility, and robust updates—we use graphs in this work.
@@ 5.2. Vector Indices in Databases
JVector (Jonathan Ellis et al., 2024) is a Cassandra-based vector index with DiskANN. They construct a DiskANN index in one shot and ingest it into the database, and support incremental insertions and deletions using FreshDiskANN. Details on how the indices are managed and updated across partitions and segments are not documented to the best of our knowledge. DataStax offers a managed product based on JVector. While their serverless offering does not provide an availability SLA, the enterprise provisioned capacity model has a 99.99% SLA. While their enterprise pricing is unspecified, their standard tier has a monthly price of \$900/month for 10 Queries/second over a million sized index. The corresponding Cosmos DB monthly cost with Autoscale is less than \$50.
Elastic (Pentreath et al., 2017) offers vector search using a segment-based indexing system. It therefore has disadvantages similar to the SingleStore-V design. Furthermore, as a managed search engine, it does not offer the same level of robustness and data consistency as as an operational database. To compare costs, consider the 10 million vector index over Wiki-Cohere data. Even with scalar quantization, we would need a machine with 15GB DRAM. We estimate using their pricing tool that the smallest such machine from Vector Search Optimized SKUs with 1.9vCPUS would costs about \$0.7/hour or about \$500/month. A manually provisioned Cosmos DB service for 500 RU/sec or 5000 RU/sec would serve 10 QPS and 100 QPS and cost about \$30\\$30\$ 30 and \$300\\$300\$ 300/month respectively.
pgvector (Andrew Kane et al., 2024) integrates HNSW and IVF-Flat indexing algorithm inside PostgreSQL and offers the inherent advantages of the PosgreSQL ecosystem. Initial versions required machines large enough to fit the entire dataset as well as the index in memory for reasonable performance. Scalar and binary quantizations have recently been added to reduce the memory requirement, but performant indexing and updates still rely on the availability of much larger memory than our system. pgvectorscale (Matvey Arye et al., 2025) similarly integrates DiskANN inside PostgreSQL and offers better performance than pgvector in limited memory settings. We leave an exhaustive benchmarking against these systems for future work. They primarily differ from our system in that they do not offer flexible schemas, multi-tenancy, and scale-out out of the box. V-Base (Zhang et al., 2023) also builds on PostgreSQL and aims to improve vector search performance. However it uses an independent buffer pool not integrated with to PostgreSQL storage engines. Substantially more work is needed to build high-availability and other features in this system.
@@ 5.3. Specialized Serverless Vector DBs
For example, Pinecone (Pinecone, 2025) offers 99.95% availability, but maintains only one data replica by default and provides limited backup-and-restore functionality applicable solely to the vector index, not the data itself. In terms of security, Pinecone supports Role-Based Access Control (RBAC) for both control and data plane operations without custom role support. Certifications include HIPAA BAA, AICPA SOC, GDPR, and ISO 27001 as of this writing.
Another offering, Zilliz (Zilliz, 2025), provides no availability SLA, lacks replication and backup-and-restore, and limits RBAC functionality to the control plane without custom roles. Its compliance certifications are limited to SOC 2 Type II, ISO/IEC 27001, and GDPR.
Turbopuffer (Eskildsen, 2024) is another serverless vector database designed for multi-tenant usage. However, it still lacks significant enterprise and robustness features such as RBAC, backup-and-restore, global availability. Turbopuffer does not publicly list query price for 10 million vector shards, however, its query complexity increases substantially with the size of the tenant. For example, cost per 1 million queries increases from \$3.58 to \$33.4 when the number of documents per name space increases from 100K to a modest 1 million(Turbopuffer, 2025).
@@ 7. Acknowledgements
We thank Philip A. Bernstein for his feedback on the work and suggestions for improving this draft.
@@ Appendix A DiskANN plots and algorithms
 return [[[closest $k$ points from $\mathcal{V}$ with label $\mathcal{B}$; $\mathcal{V}$]]]
@@ Appendix B Cosmos DB and DiskANN interoperability
To insert, the Cosmos DB IndexManager calls diskann.insert with appropriate execution context, which in turn accesses the Bw-Tree index and the Document Store. The DiskANN library written in Rust is sandwiched between IndexManager and Bw-Tree/Document Store components written in C++. The interleaving is similar for the query path, and is illustrated in Figure 15. Since all of these components are written in an asynchronous programming style with different languages and run times, this leads to an interesting challenge of implementing asynchronous callbacks across language boundary. The standard way to integrate Rust/C++ using C ABI has no native support for async calls, and public FFI libraries that support C++/Rust async inter-op are non-existent. We build an internal solution to address this, and did not observe any significant overheads from this interface.